# Learning the Mechanism of Catastrophic Forgetting: A Perspective from Gradient Similarity

### v1

@@ Appendix B Model Inference Details
Model inference was conducted in a Python (3.12.2) environment, with the main packages including transformers (4.48.0) and torch (2.2.0). To ensure experimental reproducibility, the temperature was set to 0. Model inference precision was set to FP32, and experiments were conducted on a single A800 GPU.
@@ Appendix C Model Inference Results
To identify Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$, LLMs were asked to answer the questions in the datasets. Correctly answered questions were assigned to Mastered Set $\mathcal{M}$, while incorrectly answered questions were assigned to Injection Set $\mathcal{I}$. The distribution of different sets are described in Table. 6.
@@
Model MMLU MedQA ARC-C CSQA
@@ Appendix D Model Training Details
The environment of model training was essentially the same as that for model inference. The number of epochs was set to 25. For the SGD optimizer, the learning rate included three values according to their magnitude of gradient: $2\text{e-}9$, $2\text{e-}7$, and $1\text{e-}7$ (see the code for exact configurations). For Momentum optimizer, the learning rate was consistent with SGD, and the momentum coefficient was set to 0.9. For Adam optimizer (without regularization), $\beta_{1}$ and $\beta_{2}$ were 0.9 and 0.999 respectively, and $\epsilon$ was $1\text{e-}8$. For AdamW optimizer, the regularization parameter (weight decay) was $1\text{e-}1$. Model training was performed in FP32 precision on a single NVIDIA A800 GPU.
@@ Appendix E Model Training Results
The samples in Injection Set $\mathcal{I}$ were trained for knowledge injection. After training, the number of correctly answered questions in Injection Set $\mathcal{I}$ was employed to evaluate learning, while the number of incorrectly answered questions in Mastered Set $\mathcal{M}$ was employed to evaluate forgetting. The results in Table. 7 indicates that serious catastrophic forgetting happens during knowledge injection.
@@
Model MMLU MedQA ARC-C CSQA
@@ Appendix F Evaluation on Parameter-Efficient Fine-Tuning
We evaluate LoRA as a representative parameter-efficient fine-tuning (PEFT) baseline.
All experiments are conducted on the Qwen2.5-1.5B model. For a fair comparison with standard fine-tuning (Appendix D), we adopt the same training protocol, using the SGD optimizer with a learning rate of $\eta=1\times 10^{-5}$. The LoRA configuration employs a rank of $r=16$ and a scaling factor of $\alpha=32$, with LoRA modules applied to all linear layers in the attention mechanism. The batch size and number of training epochs are identical to those used in Appendix D.
As shown in Table 8, LoRA achieves learning performance comparable to full fine-tuning, but does not alleviate catastrophic forgetting, exhibiting similarly severe forgetting across all evaluated datasets.
Table 8: Comparison of FT and LoRA under during knowledge injection. Forgetting cannot be effectively prevented when LoRA is employed.
Method MMLU MedQA ARC-C CSQA
@@
Appendix G Theoretical Derivation for More Complex Optimizers
G.1 Momentum Optimizer
$\theta$ $\leftarrow\theta-\eta m$
where $m$ performs an exponential weighted average of historical gradients for accelerated convergence. $\beta\in[0,1)$ is the momentum coefficient, used to control the decay rate of historical momentum information.
The impact of parameter updates on existing knowledge is analyzed. Similar to Eq. (2.2), we perform a Taylor expansion on loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over Mastered Set
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta m)$ (23)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta m)$
The change in loss on Mastered Set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}m}_{S^{mom}(\mathcal{M},\mathcal{I})}$ (24)
Here, $S_{mom}(\mathcal{M},\mathcal{I})$ is defined as the momentum similarity. If $S_{mom}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting. Utilizing the additivity of inner product, we decompose the global momentum similarity onto each neuron
$S^{mom}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{mom}$ (25)
where $s_{\theta_{j}}^{mom}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j}$. Based on the sign of $s_{\theta_{j}}^{mom}$, we classify the neurons into two categories:
$\theta_{CF}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})<0\right\}$ (Conflicting Neurons) (26)
$\theta_{CO}$ $=\left\{\theta_{j}\mid s_{\theta_{j}}^{mom}(\mathcal{M},\mathcal{I})\geq 0\right\}$ (Collaborative Neurons)
Similar to CNL with SGD, the parameter update rule of CNL for momentum is
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (27)
Under this rule, the loss in Master Set is
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{mom}\geq 0)\odot m\right]$ (28)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot m_{j})\right]$
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{mom}\geq 0)\cdot s_{\theta_{j}}^{mom}\right]\leq 0$
In practical settings, it is nearly impossible for all gradients to be exactly zero. Therefore we get
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (29)
G.2 Adam Optimizer (without weight decay)
The conventional update rule for the Adam optimizer is as follows:
$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (30)
$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}\quad$
$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}\quad(\text{bias correction})$
$\theta$ $\leftarrow\theta-\eta\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$
where $g_{t}=\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)$ represents the gradient vector at step $t$. $m$ and $v$ are the exponential moving average estimates of the first moment (mean) and the second raw moment (uncentered variance) of the gradients, respectively. $\beta_{1},\beta_{2}\in[0,1)$ are the exponential decay rates for the first and second moment estimates, respectively. $\hat{m}$ and $\hat{v}$ are the bias-corrected moment estimates, intended to counteract the bias caused by initializing $m,v$ to zero during the initial training phase. $\epsilon$ is a numerical stability constant added to prevent division by zero.
For notational convenience, we denote the actual update direction calculated by Adam as the vector $U_{adam}$
$U^{adam}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}$ (31)
Therefore, the conventional update rule for the Adam optimizer is
$\theta=\theta-\eta U^{adam}.$ (32)
Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adam})$ (33)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adam})$
The change in loss on the mastered set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adam}}_{S^{adam}(\mathcal{M},\mathcal{I})}$ (34)
Here, $S_{adam}(\mathcal{M},\mathcal{I})$ is defined as the Adam similarity. If $S_{adam}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
Utilizing the additivity of the inner product, we decompose the global Adam similarity onto each neuron:
$S^{adam}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adam}$ (35)
where the similarity component of the $j$-th neuron is:
$s_{\theta_{j}}^{adam}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}$ (36)
We classify the neurons into two categories:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}<0\}\quad\text{(Conflicting Neurons)}$ (37)
$\theta_{CO}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adam}\geq 0\}\quad\text{(collaborative Neurons)}$
The modified CNL parameter update formula is:
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (38)
Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{adam}\geq 0)\odot U^{adam}\right]$ (39)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adam}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adam}_{j}\right)\right]$
$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adam}\geq 0)\cdot s_{\theta_{j}}^{adam}]\leq 0$
In practical engineering,
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (40)
G.3 AdamW Optimizer
The AdamW optimizer introduces decoupled weight decay on top of Adam, separating the weight decay term from the gradient update. Its update rule is as follows:
$m$ $\leftarrow\beta_{1}m+(1-\beta_{1})g_{t}$ (41)
$v$ $\leftarrow\beta_{2}v+(1-\beta_{2})g_{t}^{2}$
$\hat{m}$ $=\frac{m}{1-\beta_{1}^{t}},\quad\hat{v}=\frac{v}{1-\beta_{2}^{t}}$
$\theta$ $\leftarrow\theta-\eta\left(\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta\right)$
where the definitions of $m,v,\hat{m},\hat{v},\beta_{1},\beta_{2},\epsilon$ are consistent with the Adam optimizer. $\lambda$ represents the Weight Decay Coefficient. Note that the update vector $U_{adamw}$ here includes not only the gradient-based adaptive step size but also the regularization term $\lambda\theta$ pointing towards the origin.
For notational convenience, we denote the actual total update direction calculated by AdamW (including weight decay) as the vector $U^{adamw}$:
$U^{adamw}=\frac{\hat{m}}{\sqrt{\hat{v}}+\epsilon}+\lambda\theta$ (42)
The rule of parameter update becomes
$\theta=\theta-\eta U^{adamw}.$ (43)
Next, we analyze the impact of parameter updates on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set:
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta U^{adamw})$ (44)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)+\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}(-\eta U^{adamw})$
The change in loss on Mastered Set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}U^{adamw}}_{S^{adamw}(\mathcal{M},\mathcal{I})}$ (45)
Here, $S_{adamw}(\mathcal{M},\mathcal{I})$ is defined as the AdamW similarity. If $S_{adamw}<0$, then $\Delta\mathcal{L}^{\mathcal{M}}>0$, leading to forgetting.
Utilizing the additivity of the inner product, we decompose the global AdamW similarity onto each neuron
$S^{adamw}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{adamw}$ (46)
where the similarity component of the $j$-th neuron is:
$s_{\theta_{j}}^{adamw}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}$ (47)
We classify the neurons into two categories:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}<0\}\quad\text{(Conflicting Neurons)}$ (48)
$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{adamw}\geq 0\}\quad\text{(collaborative Neurons)}$
The modified CNL parameter update formula is
$\theta\leftarrow\theta-\eta\left[I(S^{adamw}\geq 0)\odot U^{adamw}\right]$ (49)
Substituting the CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\cdot}\left[\mathbb{I}(\mathbf{S}^{adamw}\geq 0)\odot U^{adamw}\right]$ (50)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot\left(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot U^{adamw}_{j}\right)\right]$
$=-\eta\sum_{j=1}^{|\theta|}[I(s_{\theta_{j}}^{adamw}\geq 0)\cdot s_{\theta_{j}}^{adamw}]\leq 0$
In practical engineering
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (51)
G.4 General Optimizers
To derive a unified theoretical framework, we abstract the specific update rules of different optimizers into a generalized form. Regardless of whether the optimizer utilizes momentum, adaptive learning rates, or weight decay, its parameter update can ultimately be expressed as applying a specific update vector $F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ to the current parameters.
The general update rule is defined as:
$\theta$ $\leftarrow\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (52)
where $\eta$ is the global learning rate. Next, we analyze the impact of this general parameter update on existing knowledge. We perform a Taylor expansion on the loss function $\mathcal{L}^{\mathcal{M}}(\theta)$ over the mastered set
$\mathcal{L}^{\mathcal{M}}(\theta)$ $=\mathcal{L}^{\mathcal{M}}(\theta-\eta F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta)))$ (53)
$\approx\mathcal{L}^{\mathcal{M}}(\theta)-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$
The change in loss on the mastered set is:
$\Delta\mathcal{L}^{\mathcal{M}}\approx-\eta\underbrace{\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))}_{S^{opt}(\mathcal{M},\mathcal{I})}$ (54)
Utilizing the additivity of the inner product, we decompose the global similarity onto each neuron $j$:
$S^{opt}(\mathcal{M},\mathcal{I})=\sum_{j=1}^{|\theta|}s_{\theta_{j}}^{opt}$ (55)
where the similarity component of the $j$-th neuron is determined by the projection of the optimizer’s update scalar:
$s_{\theta_{j}}^{opt}=\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))$ (56)
Based on the sign of $s_{\theta_{j}}^{opt}$, we universally classify the neurons into two categories for any optimizer:
$\theta_{CF}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}<0\}\quad\text{(Conflicting Neurons)}$ (57)
$\theta_{CB}$ $=\{\theta_{j}\mid s_{\theta_{j}}^{opt}\geq 0\}\quad\text{(Collaborative Neurons)}$
The generalized CNL parameter update formula is:
$\theta\leftarrow\theta-\eta\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (58)
Substituting the generalized CNL update rule back into the Taylor expansion of $\Delta\mathcal{L}^{\mathcal{M}}$:
$\Delta\mathcal{L}^{\mathcal{M}}$ $\approx-\eta\nabla_{\theta}\mathcal{L}^{\mathcal{M}}(\theta)^{\top}\left[\mathbb{I}(S^{opt}\geq 0)\odot F^{opt}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$ (59)
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot(\nabla_{\theta_{j}}\mathcal{L}^{\mathcal{M}}(\theta)\cdot F^{opt}_{j}(\nabla_{\theta}\mathcal{L}^{\mathcal{I}}(\theta))\right]$
$=-\eta\sum_{j=1}^{|\theta|}\left[I(s_{\theta_{j}}^{opt}\geq 0)\cdot s_{\theta_{j}}^{opt}\right]\leq 0$
In practical engineering, we have:
$\Delta\mathcal{L}^{\mathcal{M}}<0$ (60)
Appendix H Dynamic Characteristic of CNL
The dynamic characteristics of learning and forgetting during knowledge injection are shown in Fig. 2. Although CNL freezes conflicting neurons, its training efficiency does not significantly lag behind that of conventional FT.

### v2

@@ Appendix B Model Inference Details
Model inference is conducted in a Python (3.12.2) environment, with the main packages including transformers (4.48.0) and torch (2.2.0). To ensure experimental reproducibility, the temperature is set to 0. Model inference precision is set to FP32, and experiments are conducted on a single A800 GPU.
@@ Appendix C Model Inference Results
To identify Mastered Set $\mathcal{M}$ and Injection Set $\mathcal{I}$, LLMs are asked to answer the questions in the datasets. Correctly answered questions are assigned to Mastered Set $\mathcal{M}$, while incorrectly answered questions are assigned to Injection Set $\mathcal{I}$. The distribution of different sets are described in Table. 7.
@@
MMLU MedQA ARC-C CSQA
@@ Appendix D Model Training Details
The environment of model training is essentially the same as that for model inference. The number of epochs is set to 25. For the SGD optimizer, the learning rate included three values according to their magnitude of gradient: $2\text{e-}9$, $2\text{e-}7$, and $1\text{e-}7$ (see the code for exact configurations). For Momentum optimizer, the learning rate is consistent with SGD, and the momentum coefficient is set to 0.9. For Adam optimizer (without regularization), $\beta_{1}$ and $\beta_{2}$ are 0.9 and 0.999 respectively, and $\epsilon$ is $1\text{e-}8$. For AdamW optimizer, the regularization parameter (weight decay) is $1\text{e-}1$. Model training is performed on a single NVIDIA A800 GPU.
@@ Appendix E Model Training Results
The samples in Injection Set $\mathcal{I}$ are trained for knowledge injection. After training, the number of correctly answered questions in Injection Set $\mathcal{I}$ is employed to evaluate learning, while the number of incorrectly answered questions in Mastered Set $\mathcal{M}$ is employed to evaluate forgetting. The results in Table. 8 indicates that serious catastrophic forgetting happens during knowledge injection.
@@
MMLU MedQA ARC-C CSQA
@@ Appendix F Evaluation on Parameter-Efficient Fine-Tuning
We evaluate LoRA as a representative parameter-efficient fine-tuning baseline. All experiments are conducted on the Qwen2.5-1.5B. For a fair comparison with FT, we adopt the same training protocol with Appendix D, using the SGD optimizer with a learning rate of $\eta=1\times 10^{-5}$. The LoRA configuration employs a rank of $r=16$ and a scaling factor of $\alpha=32$, with LoRA modules applied to all linear layers in the attention mechanism. The batch size and number of training epochs are identical to those used in Appendix D.
As shown in Table 9, when LoRA achieves learning performance comparable to FT, LoRA does not alleviate catastrophic forgetting, exhibiting similarly severe forgetting across all evaluated datasets.
Table 9: Comparison of FT and LoRA during knowledge injection. Forgetting cannot be effectively mitigated when LoRA is employed.
MMLU MedQA ARC-C CSQA
@@
Appendix G Extending CPL to Larger-Scale Datasets
G.1 Extending CPL to Joint Multi-Dataset Learning
Four datasets are combined to construct a more complex dataset, termed MMAC, to evaluate the performance of CPL on larger-scale datasets. Table 10 indicates that FT induced 91.7% forgetting during knowledge injection, while CPL forgets zero questions under the ideal situation that Mastered Set $\mathcal{M}$ is fully accessed.
Table 10: Performance of joint multi-dataset learning. On the larger-scale MMAC dataset, FT causes severe forgetting (91.7%), whereas CPL maintains 0.0% forgetting.
MMAC MMLU MedQA ARC-C CSQA
FT 1878 2949 625 806 665 547 316 819 272 777
CPL 1893 0 604 0 700 0 321 0 268 0
G.2 Extending CPL to Sequential Multi-Dataset Learning
The datasets are trained sequentially according to the “Order” column in Table 11. CPL consistently mitigates catastrophic forgetting across different task orders.
Table 11: Performance of sequential multi-dataset learning. AR, CS, $M$, and Me denote ARC-C, CSQA, MMLU, and MedQA, respectively. Results indicate that CPL nearly do not forget questions across all orders while maintaining effective learning, under the ideal situtation that Mastered Set $\mathcal{M}$ is fully accessed.
Order MMLU MedQA ARC-C CSQA
Learned Forgot Learned Forgot Learned Forgot Learned Forgot
AR-CS-$M$-Me 151 0 167 0 109 0 90 0
$M$-Me-AR-CS 141 1 149 0 113 0 94 0
CS-AR-Me-$M$ 150 0 148 0 106 0 98 0
Me-AR-CS-$M$ 145 0 152 0 110 0 92 0
Appendix H Dynamic Characteristic of CPL
The dynamic characteristics of learning and forgetting during knowledge injection are shown in Fig. 2. Although CPL freezes Conflicting Parameters and only updates Collaborative Parameters, the speed of knowledge injection for CPL is not noticeably slowed than FT, This indicates that Collaborative Parameters are sufficient for training, and the knowledge injection capability of CPL is not compromised by freezing Conflicting Parameters.